The potential energy U(x)of a particle moving along x-axix is given by U(x)=ax-bx2. Find the equilibrium position of particle.
1
Expert's answer
2011-09-30T10:40:56-0400
The equilibrum position is the point in which U'(x)=0.

U' = a - 2bx =
0
Hence

x = a/(2b)